Select any one public company that interests you, look up its stock symbol on the Internet, and enter this symbol into the quote lookup box under the finance link at Yahoo. Using the data available, calculate the current ratio and quick ratio for this company. Is there a difference between these two measures for this specific organization? Why or why not? If there is a difference, explain what exactly is causing it.
Answer #1

Inventory= $100,000

Cash= $50,000

Receivables= $70,000

Paybles= $60,000

Bank Over Draft= $40,000

Short term Loan= 30,000

Current Ratio= Current assets/ Current Liabilities = (100,000+50,000+70,000)/(60,000+40,000+30,000)= 1.69 times

Quick Ratio= Total Current Assets- Inventories/ Total Current Liabilities= (220,000- 100,000)/130,000= 0.92 times

In this example we got difference between Current ratio & Quick Ratio Becasue while calculating quick ratio we select all the currents except Inventories & Prepaid Expenses. We exclude Inventories because we cannot convert the invetories into cash in just few days.